OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. (8, -4) ; slope -5/6

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = -5/6 (8,-4) x1 = 8 and y1 = -4 Now just sub that into the formula and what do you get ? Don't try to solce it...just leave in the parenthesis.

OpenStudy (texaschic101):

typo.....don't try to solve it